# Nonadiabatic corrections to electric quadrupole transition rates in H2

### v1

@@ 7 Summary
With the improved quadrupole moment curve, one may recalculate the whole E2 spectrum of H2, but the ultimate way is the direct nonadiabatic calculation currently being pursued by Komasa [21]. In addition, our results can be extended to heteronuclear molecules (such as HD) without major modifications. Moreover, analogous formulas for nonadiabatic corrections can be derived for the electric polarizability and QED corrections to transition energies.
@@ Appendix A Alternative Calculation of $Q_{2}$
Summarizing, one calculates $Q_{2}$ using Eq. (98), with $\vec{v}$ obtained from Eq. (69), $\partial_{R}\vec{v}$ from Eq. (72), $\delta\vec{v}$ from Eq. (101), $\partial_{R}\delta\vec{v}$ from Eq. (A), $\partial_{R}\mathcal{E}_{\mathrm{el}}$ from (107), and with $\partial_{R}\mathcal{E}_{\mathrm{el}}$ from (109).

### v2

@@ 7 Summary
With the improved quadrupole moment curve, one may recalculate all transition rates and lifetimes of H2 (and D2, T2), which will be the subject of future work. The ultimate way, however, is the direct nonadiabatic calculation currently being pursued by Komasa [26]. In addition, our results can be extended to heteronuclear molecules (such as HD) without major modifications. Moreover, we note that an analogous method can be used to derive formulas for nonadiabatic corrections for the electric polarizability and QED corrections to transition energies.
@@ Acknowledgements
Disclosure statement
No potential conflict of interest was reported by the authors.
@@ Appendix A Alternative Calculation of $Q_{2}$
Summarizing, one calculates $Q_{2}$ using Eq. (98), with $\vec{v}$ obtained from Eq. (69), $\partial_{R}\vec{v}$ from Eq. (72), $\delta\vec{v}$ from Eq. (101), $\partial_{R}\delta\vec{v}$ from Eq. (A), $\partial_{R}\mathcal{E}_{\mathrm{el}}$ from (107), and with $\partial_{R}\delta\mathcal{E}_{\mathrm{el}}$ from (109).